OpenStudy (anonymous):

Match Term Definition negative sixteen divided by negative four A) negative eight negative twenty−four divided by six B) eight negative twenty−four divided by negative three C) four sixteen divided by negative two D) negative four

OpenStudy (anonymous):

@piglet9

OpenStudy (anonymous):

@12rabbits

OpenStudy (anonymous):

-16/-4 = 4 = C -24/6 = -4 = D -24/-3 = 8 = B 12/-2 = -8 = A

OpenStudy (anonymous):

wait the last one should be 16 /2 not 12/2

OpenStudy (anonymous):

but it doesn't matter lol

OpenStudy (anonymous):

yeah i guessed that too lol

OpenStudy (mathstudent55):

The answer is still correct despite the fact that the 16 was mistakenly written as 12. 16/(-2) = -8. Answer is A.

OpenStudy (anonymous):

we know